in 3 years, alex will be 3 times as old as prency. a year ago, alex was 7 times as old as prency. how old are they now?
Answer by mananth(16946)

prency = x
alex = y
In 3 years
y+3=3(x+3)
y+3=3x+9
y-3x=6..........1
1 year ago
y-1=7(x-1)
y-1=7x-7
y-7x=-6............2
-3 x + 1 y = 6 .............1
-7 x + 1 y = -6 .............2

multiply (1)by 1
Multiply (2) by -1
-3 x + 1 y = 6
7 x + -1 y = 6
Add the two equations
4 x = 12
/ 4
x = 3
plug value of x in (1)
-3 x + 1 y = 6
-9 + 1 y = 6
1 y = 6 + 9
1 y = 15
y = 15